OpenStudy (anonymous):

Can someone walk me through the steps of this quadratic formula? -x^2-3+5=0 Thank you!

hartnn (hartnn):

yes, i assume you have the formula with you... could you find the values of a,b,c in it ?

hartnn (hartnn):

and i also assume its \(-x^2-3x+5=0\)

OpenStudy (anonymous):

If it is like the first formula, I would start with -x^2+3-5=0 I think

hartnn (hartnn):

\(\large x= \dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (anonymous):

I am unfamiliar with using variables with this formula.

hartnn (hartnn):

compare your equation with \(ax^2+bx+c=0\) what you get as values of a,b,c ?

OpenStudy (anonymous):

Standard form of a parabola \[ax ^{2}+bx+c\]

OpenStudy (anonymous):

BTW so you mean -x^2-3x+5 instead of -x^2-3+5?

OpenStudy (anonymous):

Do** not so

OpenStudy (anonymous):

-x^2-3+5=0 is the formula is that what you meant?

OpenStudy (anonymous):

okay well that simplifies to -x^2+2 and to find what a b and c are we are going to compare -x^2+2 to \[ax ^{2}+bx+c\]

OpenStudy (anonymous):

Ok, I am writing the steps so that I can complete the other problems on my own.

OpenStudy (anonymous):

I'm going to start over and write in steps for you!

OpenStudy (anonymous):

That would be great!

OpenStudy (anonymous):

Step 1: write quadratic equation! \[\frac{-b \pm \sqrt{b ^{2}-4(a)(c)}}{ 2a }\]

OpenStudy (anonymous):

Step two: compare your function \[-x ^{2}+2\] in this case with \[a x^{2}+bx+c\]

OpenStudy (anonymous):

Step 3: find a,b, and c \[-x ^{2}+2\] has an "a" of -1, a "b" of 0, and a "c" of 2 Tell me if you understand where I got that from.

OpenStudy (anonymous):

You are following the formula that you wrote first ax^2+bx+c

OpenStudy (anonymous):

Right! Okay let's move on

OpenStudy (anonymous):

Step 4: plug your values of a, b, and c into the quadratic equation. a=-1 b=0 c=2, and plug them into the quadratic equation of \[\frac{ -b \pm \sqrt{b ^{2}-4(a)(c)} }{ 2a }\] then we get \[\frac{ 0\pm \sqrt{0-4(-1)(2)} }{ 2(-1) }\]

OpenStudy (anonymous):

Step 5: simplify \[\frac{ 0\pm \sqrt{0-4(-1)(2)} }{ 2(-1) } = \frac{ 0\pm \sqrt{8} }{ -2 }\]

OpenStudy (anonymous):

Step 6: evaluate \[0+ -\frac{ \sqrt{8} }{ 2 } = -\frac{ \sqrt{8} }{ 2 } and 0 - -\frac{ \sqrt{8} }{ 2 } = \frac{ \sqrt{8} }{ 2 }\]

OpenStudy (anonymous):

Or you could use decimals sqrt(8)= 2.83 (2.83)/-2= -1.42 then we would have \[0+(-1.42)= -1.42 \] and \[0-(-1.42)= 1.42\]

OpenStudy (anonymous):

Step 7: understand what we just found! We found the points where this parabola crosses the x-axis, also known as the x-intercepts!

OpenStudy (anonymous):

I will use step 6 because I understood every step that you wrote. Thank you for Step 7 because I could never make this connection before now. I will work these problems in the math lab now. Thank you and here is your medal!

OpenStudy (anonymous):

Thanks man good job. If you wanna see what we just did go to https://www.desmos.com/calculator and type in -x^2+2, then put your mouse over the points where the graph crosses the x axis!

OpenStudy (anonymous):

I will do that, you have just gotten a new fan! Thank you!

OpenStudy (anonymous):

Glad to help!

OpenStudy (anonymous):

It will take me a lot longer to solve but I now have the complete formula! Thanks again!

OpenStudy (anonymous):

:)